103. The most recent financial statements for Moose Tours, Inc. follow. Sales for 2013 are projected to grow by 16 percent. Interest expense will remain constant; the tax rate and dividend payout rate will also remain constant. Costs, other expenses, current assets, and accounts payable increase spontancously will sales. If the firm is operating at full capacity and no new debt or sales? equity is issued, how much external financing is needed to support the 16 percent growth rate in 2012 Income Statement Sales Costs Other expenses Earnings before interest and taxes Interest paid Taxable income Taxes (34%) Net income $910,000 709,000 $190,000 20.900 $169,100 S111.606 Dividends Addition to retained earnings $44.642 $66.964 Balance Sheet as of December 31, 2012 Assets Liabilities and Owners Equity Current assets Cash Accounts receivable Inventory Current liabilities Notes payable Long-term debt $24,000 42,000 Accounts payable $66,000 10,000 $76,000 $142,000 Total Total Fixed assets $142,000 Owners equity Net plant and equipment 367000 Common stock and paid-in surplus $23,000 268,000 $291,000 Retained earnings Total Total liabilities and owners equity Total assets $509,000 $509,000 A. $-10,246 B. -$8,122 C. -$6,708 D. $2,407 E. $3,309
Answer #1

Step 1: Prepare Proforma Income Statement

The proforma income statement is prepared as below:

Proforma Income Statement
Sale [910,000*(1+16%)] 1,055,600
Less Costs [709,000*(1+16%)] 822,440
Other Expenses [11,000*(1+16%)] 12,760
EBIT 220,400
Less Interest Paid 20,900
Taxable Income 199,500
Less Taxes (199,500*34%) 67,830
Net Income $131,670

______

Step 2: Calculate Estimated  Dividend, Addition to Retained Earnings and Ending Retained Earnings

The value of revised dividends, addition to retained earnings and ending retained earnings is arrived as follows:

Estimated Dividend for 2013 = Dividends for 2012/Net Income for 2012*Estimated Net Income for 2013 = 44,642/111,606*131,670 = $52,668

Addition to Retained Earnings = Estimated Net Income for 2013 - Estimated Dividend for 2012 = 131,670 - 52,668 = $79,002

Ending Retained Earnings = Retained Earnings as Per 2012 Balance Sheet + Addition to Retained Earnings = 268,000 + 79,002 = $347,002

______

Step 3: Prepare Proforma Balance Sheet

The proforma balance sheet is prepared as below:

Proforma Balance Sheet
Assets Liabilities and Owner's Equity
Current Assets Current Liabilities
Cash [24,000*(1+16%)] 27,840 Accounts Payable [66,000*(1+16%)] 76,560
Accounts Receivable [42,000*(1+16%)] 48,720 Notes Payable 10,000
Inventory [76,000*(1+16%)] 88,160 Total 86,560
Total 164,720 Long-Term Debt 142,000
Fixed Assets Owner's Equity
Net Plant and Equipment [367,000*(1+16%)] 425,720 Common Stock and Paid-in Surplus 23,000
Retained Earnings 347,002
Total 370,002
Total Assets $590,440 Total Liabilities and Stock Owner's Equity $598,562

______

Step 4: Calculate External Financing Needed

The value of external financing needed is determined as below:

External Financing Needed = Total Assets as Per Proforma Balance Sheet - Total Liabilities and Stock Owner's Equity as Per Proforma Balance Sheet = 590,440 - 598,562 = -$8,122 (which is Option B)